OpenStudy (anonymous):

Ryan flew from Wiley Post to Ponca City and back. Ryan maintained an average rate of 450 mph going to Ponca City and an average rate of 400 mph returning to Wiley Post. If the actual flying time for the roundtrip was 1 hour, about how far is it from Wiley Post to Ponca City? Round to nearest mile. A. 0.47 miles B. 212 miles C. 239 miles D. 3600 miles

OpenStudy (anonymous):

450(t) = 400(1-t) 850t = 400 => t = 400/850 => t = 0.47 hours. => distance = 450*0.47 = 212 miles.

OpenStudy (anonymous):

what gt did there was used distance = rate * time 450 was the rate there and he used t as the time it took to get there, so that means the time it took to get back was 1 - t because flight there and flight back added up to 1 hour.

OpenStudy (anonymous):

when he solved for t, that was the time it took to fly there, and distance = rate * time so D = 450 * .47 = 212 miles